# Locating Superspreaders in SI Cascades

### 05 Mar 2024 v1

@@ 6 Organizing second deriviative stuff
$|{\mathcal{Q}}|\leq nd^{2}+mnd\leq 2mnd^{2}.$ (6.1)
The first inequality can be explained as follows. If $v\in{\mathcal{I}}[0,t]$ and $v$ has degree at most $d$, the number of paths $P\in{\mathcal{Q}}$ which begin at $v$ is at most $d^{2}$. As there are at most $n$ vertices in ${\mathcal{I}}[0,t]$ of degree at most $d$, the total number of $P\in{\mathcal{Q}}$ starting at such a vertex is $nd^{2}$. On the other hand, if $v\in\mathsf{highdeg}(G)\cap{\mathcal{I}}[0,t]$, then the total number of paths of length 2 beginning at $v$ is at most $nd$, since $\deg(v)\leq n$. As there are $m$ vertices in $\mathsf{highdeg}(G)$, the number of paths in ${\mathcal{Q}}$ for this case is at most $mnd$. Summing both together leads to the claimed bound.
@@
($\operatorname{\mathbb{P}}\left(\{|{\mathcal{I}}_{2}|\geq 1\}\cap{\mathcal{A}}\cap{\mathcal{B}}_{t}|{\mathcal{F}}_{t}\right)$ $\leq\operatorname{\mathbb{P}}\left(\exists P\in{\mathcal{Q}}:\mathsf{weight}(P)\leq\delta|{\mathcal{F}}_{t}\right)$
$\leq\sum_{P\in{\mathcal{Q}}}\operatorname{\mathbb{P}}(\mathsf{weight}(P)\leq\delta|{\mathcal{F}}_{t})$
$\stackrel{{\scriptstyle(a)}}{{\leq}}|{\mathcal{Q}}|\lambda^{2}\delta^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}2mnd^{2}\lambda^{2}\delta^{2}=n^{-(2\beta-1)+o(1)}.$
@@
($\operatorname{\mathbb{P}}\left(\{|{\mathcal{I}}_{2}|\geq 1\}\cap{\mathcal{A}}|{\mathcal{F}}_{t}\right)$ ($\leq\operatorname{\mathbb{P}}\left(\{|{\mathcal{I}}_{2}|\geq 1\}\cap{\mathcal{A}}\cap{\mathcal{B}}_{t}|{\mathcal{F}}_{t}\right)+\operatorname{\mathbb{P}}({\mathcal{A}}\cap{\mathcal{B}}_{t}^{c}|{\mathcal{F}}_{t})$
@@
We now put everything together to prove the desired result.
@@
For any $t\geq 0$, Lemmas 5.2 and 6.5 imply that
($\operatorname{\mathbb{P}}\left(\left.\left\{\left||{\mathcal{I}}(t,t+\delta]-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right|{\mathcal{F}}_{t}\right)\\ \leq\operatorname{\mathbb{P}}\left(\left.\left\{\left||{\mathcal{I}}_{1}|-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right|{\mathcal{F}}_{t}\right)+\operatorname{\mathbb{P}}(\{|{\mathcal{I}}_{2}|\geq 1\}\cap{\mathcal{A}}|{\mathcal{F}}_{t})\leq n^{-\left(\beta-\frac{1}{2}\right)+o(1)}.$
Through identical reasoning as the proof of Lemma [reF], it follows that
@@
($\operatorname{\mathbb{P}}\left(\left.\left\{\left|\left|{\mathcal{I}}\left(T(v),T(v)+\frac{1}{n^{\beta}}\right]\right|-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right|{\mathcal{F}}_{T(v)}\right)\leq n^{-\left(\beta-\frac{1}{2}\right)+o(1)},$
which holds since $T(v)<\infty$ almost surely. Finally, the desired result follows from taking an expectation with respect to ${\mathcal{F}}_{T(v)}$. ∎
7 Algorithm that works for all $\alpha\in(0,1)$
[high-level idea] For each vertex, look at the number of consistent infections that occur in 1 timestep after it is infected. If we look at the set of vertices that have at least $D/2$ vertices in their consistent infection set, this will be a subgraph of size $n^{\alpha+\epsilon}$. We can then do the same second-derivative test to find the super-spreaders.
[update: this doesn’t work!!] At least not for general graphs, due to the many possible paths that can lead to the shortest one. So let’s just throw in the proof for trees, and emphasize the difficulties. General graphs can be an open problem.
[formal results]
[a few results on comparing gamma variables]
[possible notation:] When there is a single high-degree vertex, denote it by $v_{\mathrm{HD}}$.
Lemma 7.1.
Suppose that $X\sim\mathrm{Gamma}(a,\lambda)$ and $Y\sim\mathrm{Gamma}(b,\lambda)$ are independent with $1001\leq a\leq b/$ 100. Then
$\operatorname{\mathbb{P}}(Y\leq X-1)\leq e^{\lambda}\left(\frac{2}{3}\right)^{a+b}.$
Proof.
For any $\theta\in(0,\lambda)$, it holds that
$\operatorname{\mathbb{P}}(Y\leq X-1)=\operatorname{\mathbb{P}}(X-Y\geq-1)\leq e^{\theta}\operatorname{\mathbb{E}}[e^{\theta(X-Y)}]\leq e^{\lambda}\left(1-\frac{\theta}{\lambda}\right)^{-a}\left(1+\frac{\theta}{\lambda}\right)^{-b}$
Setting $\theta:=\lambda(b-a)/(a+b)$, we arrive at the inequality
$\operatorname{\mathbb{P}}(Y\leq X-1)\leq\exp\left\{\lambda-(a+b)\left(\log 2-H\left(\frac{a}{a+b}\right)\right)\right\},$ (7.1)
where we define the binary entropy function $H(p):=-p\log p-(1-p)\log(1-p)$. Since $b/a\geq 100$, we have that
$H\left(\frac{a}{a+b}\right)\leq H\left(\frac{1}{101}\right)\leq\log\left(\frac{4}{3}\right),$
and it follows that
$\log 2-H\left(\frac{a}{a+b}\right)\geq\log\left(\frac{3}{2}\right).$
Substituting this inequality into (7.1) implies the claimed result. ∎
Lemma 7.2.
Let $X\sim\mathrm{Gamma}(a,\lambda)$ with $a\geq 1$. For $a$ sufficiently large, it holds that
$\sup_{x>0}\operatorname{\mathbb{P}}\left(|X-x|\leq 1\right)\leq\frac{\lambda}{\sqrt{a}}.$
Proof.
Letting $f(x)$ be the density function for $X$, define $f_{\sup}:=\sup_{x>0}f(x)$ to be its maximum value. Since we can bound
[…]
it suffices to characterize $f_{\sup}$. To this end, we have that
$f_{\sup}$ $:=\sup_{x>0}\frac{\lambda^{a}}{\Gamma(a)}x^{a-1}e^{-\lambda x}\stackrel{{\scriptstyle(a)}}{{=}}\frac{\lambda}{\Gamma(a)}\left(a-1\right)^{a-1}e^{-(a-1)}\stackrel{{\scriptstyle(b)}}{{\sim}}\frac{\lambda}{\sqrt{2\pi a}}.$
Above, $(a)$ uses the fact that the mode of a $\mathrm{Gamma}(a,\lambda)$ random variable is $(a-1)/\lambda$, and $(b)$ is due to Stirling’s approximation. Substituting this asymptotic expression into (7.2) shows that for $a$ sufficiently large,
$\operatorname{\mathbb{P}}(|X-x|\leq 1)\lesssim\frac{2\lambda}{\sqrt{2\pi a}}\leq\frac{\lambda}{\sqrt{a}}.$
[narrowing down the search space]
[need to set the right value of $R$ and $K$]
Lemma 7.3.
It holds with probability $1-o(1)$ that for all $v\in V$,
$\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]\subset{\mathcal{N}}(v,R).$
Proof.
As a shorthand, we set $R:=\epsilon\log(n)/\log(d)$ in this proof. Fix vertices $u,v\in V$ with $\mathsf{dist}_{G}(u,v)\geq R$ as well as $i\in[K]$. Let $P_{i}$ and $Q_{i}$ be the unique paths from the source vertex $s_{i}$ to $u_{i}$ and $v_{i}$, respectively. If $P_{i}$ and $Q_{i}$ are disjoint, then it holds that $T_{i}(u)=\mathsf{weight}_{i}(P_{i})$ and $T_{i}(v)=\mathsf{weight}_{i}(Q_{i})$ are independent, with $\mathsf{weight}_{i}(P_{i})\sim\mathrm{Gamma}(|P_{i}|,\lambda)$ and $\mathsf{weight}_{i}(Q_{i})\sim\mathrm{Gamma}(|Q_{i}|,\lambda)$. As the graph is a tree, it holds that $|P_{i}|+|Q_{i}|=\mathsf{dist}_{G}(u,v)\geq R$, hence we have that either $|P_{i}|\geq R/2$ or $|Q_{i}|\geq R/2$. Lemma 7.2 therefore implies
$\operatorname{\mathbb{P}}\left(|T_{i}(u)-T_{i}(v)|\leq 1\right)\leq\frac{\lambda}{\sqrt{R/2}}.$ (7.3)
In the case where $P_{i}$ and $Q_{i}$ are not disjoint, let $w$ be the last common element of both paths. We then define $P_{i}^{\prime}$ and $Q_{i}^{\prime}$ to be the portion of $P_{i}$ and $Q_{i}$, respectively, that starts at $w$ instead of $s_{i}$. By definition, $P_{i}^{\prime}$ and $Q_{i}^{\prime}$ are disjoint, and we have that $T_{i}(u)=T_{i}(w)+\mathsf{weight}_{i}(P_{i}^{\prime})$ and $T_{i}(v)=T_{i}(w)+\mathsf{weight}_{i}(Q_{i}^{\prime})$. Moreover, $\mathsf{weight}_{i}(P_{i}^{\prime})\sim\mathrm{Gamma}(|P_{i}^{\prime}|,\lambda)$ and $\mathsf{weight}(Q_{i}^{\prime})\sim\mathrm{Gamma}(|Q_{i}^{\prime}|,\lambda)$, are independent, and as before, $|P_{i}^{\prime}|+|Q_{i}^{\prime}|=\mathsf{dist}_{G}(u,v)\geq R$. By the same logic as the case where $P_{i},Q_{i}$ are disjoint, we see that (7.3) holds in this case as well.
From (7.3), it is readily seen that for $K$ satisfying [eqref], it holds that
$\operatorname{\mathbb{P}}\left(|T_{i}(u)-T_{i}(v)|\leq 1,\forall i\in[K]\right)=o\left(\frac{1}{n^{2}}\right).$ (7.4)
Finally, to prove the claim, we can bound
$\operatorname{\mathbb{P}}\left(\exists v\in V:\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]\not\subset{\mathcal{N}}(v,R)\right)$ $=\operatorname{\mathbb{P}}\left(\sum_{v\in V}\left|\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]\setminus{\mathcal{N}}(v,R)\right|\geq 1\right)$
@@ 7 Algorithm that works for all $\alpha\in(0,1)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{v\in V}\operatorname{\mathbb{E}}\left[\left|\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]\setminus{\mathcal{N}}(v,R)\right|\right]$
$=\sum_{v\in V}\sum_{u\in V\setminus{\mathcal{N}}(v,R)}\operatorname{\mathbb{P}}\left(u\in\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]\right)$
$\leq\sum_{v\in V}\sum_{u\in V\setminus{\mathcal{N}}(v,R)}\operatorname{\mathbb{P}}\left(\left|T_{i}(u)-T_{i}(v)\right|\leq 1,\forall i\in[K]\right)$
=$\stackrel{{\scriptstyle(b)}}{{=}}o(1).$
Above, $(a)$ is due to Markov’s inequality, and $(b)$ follows from (7.4). ∎
Define the set
$\widehat{U}:=\left\{v\in V:\left|\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]\right|>n^{\alpha/2}\right\}.$
[set $R$ below]
Lemma 7.4.
With probability $1-o(1)$, the following hold:
  1. 1.
    $\widehat{U}\subset{\mathcal{N}}(v_{\mathrm{HD}},R)$.
  2. 2.
    $v_{\mathrm{HD}}\in\widehat{U}$.
  3. 3.
    $|\widehat{U}\cap{\mathcal{N}}(v_{\mathrm{HD}})|\geq\frac{1}{4}(1-e^{-\lambda/2})^{K}|{\mathcal{N}}(v_{\mathrm{HD}})|$.
Proof.
Suppose that $u\notin{\mathcal{N}}(v_{\mathrm{HD}},R)$, which is equivalent to $v_{\mathrm{HD}}\notin{\mathcal{N}}(u,R)$. By Lemma 7.3, it holds with probability $1-o(1)$ that for all such $u$,
$\left|\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(u),T_{i}(u)+1]\right|\leq|{\mathcal{N}}(u,R)|\leq d^{R}=o(n^{\alpha/2}).$ (7.5)
[explain upper bound on $|{\mathcal{N}}(u,R)|$] In particular, (7.5) implies that $\widehat{U}$ is contained in ${\mathcal{N}}(v_{\mathrm{HD}},R)$ with probability $1-o(1)$, proving Item #1.
Next, we turn to the proof of Item #2. For each $i\in[K]$, we let $v_{i}$ be the penultimate vertex on the unique path from $s_{i}$ to $v_{\mathrm{HD}}$ if $s_{i}\neq v$, otherwise we set $v_{i}:=v_{\mathrm{HD}}$. Notice that for $u\in{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{i}\}$, it holds that
$T_{i}(u)=T_{i}(v_{\mathrm{HD}})+E_{i}(u),\qquad u\in{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{i}\},i\in[K],$
where the $E_{i}(u)$’s are i.i.d. $\mathrm{Exp}(\lambda)$ random variables. Notice that for $u\in{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{1},\ldots,v_{K}\}$,
$E_{i}(u)\leq 1,\forall i\in[K]\Rightarrow u\in\bigcap_{i=1}^{K}{\mathcal{I}}_{i}[T_{i}(v_{\mathrm{HD}}),T_{i}(v_{\mathrm{HD}})+1].$
Since $\operatorname{\mathbb{P}}(\max_{i\in[K]}E_{i}(u)\leq 1)=(1-e^{-\lambda})^{K}$, the $E_{i}(u)$’s are i.i.d., and $|{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{1},\ldots,v_{K}\}|\geq n^{\alpha}-K\geq n^{\alpha}/2$, it follows that $|\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]|$ stochastically bounds a $\mathrm{Bin}(n^{\alpha}/2,(1-e^{-\lambda})^{K})$ random variable. It now follows from Bernstein’s inequality that
$\operatorname{\mathbb{P}}\left(\left|\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]\right|\leq\frac{1}{4}(1-e^{-\lambda})^{K}n^{\alpha}\right)=o(1).$
Item #2 immediately follows, since $(1-e^{-\lambda})^{K}n^{\alpha}/4\geq n^{\alpha/2}$ for $n$ sufficiently large.
We now prove Item #[ref]. Let us define the sets
$S_{1}$ :=$:=\{u\in{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{1},\ldots,v_{K}\}:E_{i}(u)\in[0,1/2],\forall i\in[K]\}$
$S_{2}$ :=$:=\{u\in{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{1},\ldots,v_{K}\}:E_{i}(u)\in[1/2,1],\forall i\in[K]\}.$
For $u\in S_{1}$ and $w\in S_{2}$, it holds for all $i\in[K]$ that $T_{i}(w)\in[T_{i}(u),T_{i}(u)+1]$, hence
$S_{2}\subset\bigcap_{i=1}^{K}{\mathcal{I}}_{i}[T_{i}(u),T_{i}(u)+1],\qquad,\forall u\in S_{1}.$
In particular, [eqref] implies that $|\widehat{U}\cap{\mathcal{N}}(v_{\mathrm{HD}})|\geq|S_{1}|$ if $|S_{2}|\geq n^{\alpha/2}$. We therefore proceed by establishing lower bounds for $|S_{1}|$ and $|S_{2}|$. For $u\in{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{1},\ldots,v_{K}\}$, it holds that
$\operatorname{\mathbb{P}}(u\in S_{1})$ $=\prod_{i=1}^{K}\operatorname{\mathbb{P}}(E_{i}(u)\in[0,1/2])=(1-e^{-\lambda/2})^{K}$
$\operatorname{\mathbb{P}}(u\in S_{2})$ $=\prod_{i=1}^{K}\operatorname{\mathbb{P}}(E_{i}(u)\in[1/2,1])=\left(e^{-\lambda/2}(1-e^{-\lambda/2})\right)^{K}.$
As the $E_{i}(u)$’s are independent, we may apply Bernstein’s inequality to obtain that
$|S_{1}|$ $\geq\frac{1}{2}(1-e^{-\lambda/2})^{K}|{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{1},\ldots,v_{K}\}|\geq\frac{1}{4}(1-e^{-\lambda/2})^{K}|{\mathcal{N}}(v_{\mathrm{HD}})|$
$|S_{2}|$ $\geq\frac{1}{2}\left(e^{-\lambda/2}(1-e^{-\lambda/2})\right)^{K}|{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{1},\ldots,v_{K}\}|\geq\frac{1}{4}\left(e^{-\lambda/2}(1-e^{-\lambda/2})\right)^{K}|{\mathcal{N}}(v_{\mathrm{HD}})|,$
with probability $1-o(1)$. In the final inequality on both lines, we have used that |$|\{v_{1},\ldots,v_{K}\}|\leq K\leq|{\mathcal{N}}(v_{\mathrm{HD}})|/2$. In particular, $|S_{2}|\geq n^{\alpha/2}$ with probability $1-o(1)$, so it follows that
$|\widehat{U}\cap{\mathcal{N}}(v_{\mathrm{HD}})|\geq|S_{1}|\geq\frac{1}{4}(1-e^{-\lambda/2})^{K}|{\mathcal{N}}(v_{\mathrm{HD}})|,$
with probability $1-o(1)$. ∎
7.1 Proof of restricted cut thing
[next result: for a given set $S$, $|{\mathcal{I}}[T(v),T(v)+\delta]\cap S|$ concentrates around $|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,T(v)]\cap S)|$.]
Let $S\subset V$, fix $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Recall the definitions of ${\mathcal{I}}_{1}$ and ${\mathcal{I}}_{2}$ from Section [ref] (see [eqrefs]); the objective of this section is to study $|{\mathcal{I}}_{1}\cap S|$ and $|{\mathcal{I}}_{2}\cap S|$. The techniques needed for doing so are largely the same as those in Section [ref], and we emphasize the key differences here.
We begin by studying the size of ${\mathcal{I}}_{1}\cap S$.
Lemma 7.5.
Let $\gamma>\max\{\alpha/2,1-2\beta\}$ and let $t\geq 0$. Then it holds with probability $1-o(1/n)$ that
@@ 7.1 Proof of restricted cut thing
$\operatorname{\mathbb{P}}\left(\left.\left||{\mathcal{I}}_{1}\cap\widehat{U}|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t]\cap\widehat{U})|\right|\geq n^{\gamma}\right|{\mathcal{F}}_{t}\right)=o\left(\frac{1}{n}\right).$
Remark 7.6.
The event ${\mathcal{A}}$ is not needed here, since in trees, there is a single infection path from the source to any given vertex.
Proof.
Condition on ${\mathcal{F}}_{t}$ as well as $\widehat{U}$. Since $G$ is a tree, we have for each $v\in\partial{\mathcal{I}}(t)$ that $v$ has exactly one infected neighbor at time $t$. Hence $T(v)-t\sim\mathrm{Exp}(\lambda)$ conditioned on ${\mathcal{F}}_{t}$, so
$\operatorname{\mathbb{P}}(v\in{\mathcal{I}}_{1}|{\mathcal{F}}_{t})=\operatorname{\mathbb{P}}(T(v)-t\in[0,n^{-\beta}]|{\mathcal{F}}_{t})=1-e^{-\lambda/n^{\beta}}.$
By the conditional independence of infection times in $\partial{\mathcal{I}}(t)$ with respect to ${\mathcal{F}}_{t}$ and $\widehat{U}$, we have that $|{\mathcal{I}}_{1}|\sim\operatorname{Bin}(|\partial{\mathcal{I}}(t)|,1-e^{-\lambda/n^{\beta}})$.
By the definition of ${\mathcal{I}}_{1}$, we have that ${\mathcal{I}}_{1}\cap S\subset\partial{\mathcal{I}}(t)\cap\{v\in V:{\mathcal{N}}(v)\cap\widehat{U}\neq\emptyset\}$. Observe further that each $v\in\partial{\mathcal{I}}(t)$ has exactly one neighbor in ${\mathcal{I}}[0,t]$. Indeed, if there were at least two infected neighbors of $v$, there must exist two distinct paths from the cascade source $s$ to $v$, which is not possible under our assumption that the underlying graph is a tree. Hence it holds for $v\in\partial{\mathcal{I}}(t)$ that
$\operatorname{\mathbb{P}}(v\in{\mathcal{I}}_{1}|{\mathcal{F}}_{t})=1-e^{-\lambda\delta}.$
Following the proof of Lemma [ref], we have the representation
$|{\mathcal{I}}_{1}\cap\widehat{U}|\stackrel{{\scriptstyle d}}{{=}}\sum_{v\in\partial{\mathcal{I}}(t):{\mathcal{N}}(v)\cap\widehat{U}\neq\emptyset}X(v),$
where the $X(v)$’s are independent with $X(v)\sim\mathrm{Bern}\left(1-e^{-\lambda\delta}\right)$. Analogously to [eqref], Bernstein’s inequality implies that for any $x\geq 0$,
$\operatorname{\mathbb{P}}\left(\left.\left||{\mathcal{I}}_{1}\cap\widehat{U}|-\operatorname{\mathbb{E}}[|{\mathcal{I}}_{1}\cap\widehat{U}||{\mathcal{F}}_{t},\widehat{U}]\right|\geq x\right|{\mathcal{F}}_{t},\widehat{U}\right)\leq 2\exp\left(-\frac{-x^{2}/2}{|\partial{\mathcal{I}}(t)\cap\{v:{\mathcal{N}}(v)\cap\widehat{U}\neq\emptyset\}|\left(1-e^{-\lambda\delta}\right)+x/3}\right).$
To simplify the bound further, we make a number of observations. First, since the graph $G$ is a tree, we have that
$|\partial{\mathcal{I}}(t)\cap\{v\in V:{\mathcal{N}}(v)\cap\widehat{U}\neq\emptyset\}|=|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t]\cap\widehat{U})|.$
Moreover, we can bound
$|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t]\cap\widehat{U})|\leq|\operatorname{\mathsf{cut}}(\widehat{U})|\leq\sum_{v\in\widehat{U}}|{\mathcal{N}}(v)|\leq|{\mathcal{N}}(v_{\mathrm{HD}})|+d|\widehat{U}|\leq n^{\alpha+\epsilon}.$
[explain.] As a consequence of [eqref], we can characterize the conditional expectation of $|{\mathcal{I}}_{1}\cap\widehat{U}|$ as follows.
$\left|\operatorname{\mathbb{E}}[|{\mathcal{I}}_{1}\cap\widehat{U}||{\mathcal{F}}_{t},\widehat{U}]-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t]\cap\widehat{U})|\right|\leq\frac{\lambda^{2}}{n^{2\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t]\cap\widehat{U})|\leq\lambda^{2}n^{\alpha-2\beta+\epsilon}\leq\frac{1}{2}n^{\gamma}.$
[explain] Setting $x=n^{\gamma}/2$ in [eqref] as well as using the inequalities [eqref] and $1-e^{-\lambda\delta}\leq\lambda\delta$ proves the claim. ∎
[Next result: bounding $|{\mathcal{I}}_{2}\cap\widehat{U}|$]
8 Lower bounds for $\alpha\in(0,1/2)$
Theorem 8.1.
Fix $\epsilon>0$ and $\alpha\in(0,1/2)$. Let $D=n^{\alpha+o(1)}$ and let ${\mathcal{G}}$ be the class of all graphs where all vertices have degree at most $\log^{2}n$, except for at most one vertex with degree at least $D$. If
$K\leq\left(\frac{1-2\alpha-\epsilon}{5}\right)\frac{\log n}{\log\log n},$
then
@@ 8 Lower bounds for $\alpha\in(0,1/2)$
$\inf_{\widehat{v}}\max_{G\in{\mathcal{G}}}\operatorname{\mathbb{P}}_{G}\left(\widehat{v}(\mathbf{T})\neq\mathsf{highdeg}(G)\right)=1-o(1).$

### 06 Mar 2024 v2

@@ 3 Preliminaries: Cascade dynamics and infection times
$|{\mathcal{Q}}|\leq nd^{2}+mnd=n^{1+o(1)}.$ (5.7)
The first inequality can be explained as follows. If $v\in{\mathcal{I}}(t)$ and $v$ has degree at most $d$, the number of paths $P\in{\mathcal{Q}}$ which begin at $v$ is at most $d^{2}$. As there are at most $n$ vertices in ${\mathcal{I}}(t)$ of degree at most $d$, the total number of $P\in{\mathcal{Q}}$ starting at such a vertex is $nd^{2}$. On the other hand, if $v\in\mathsf{highdeg}(G)\cap{\mathcal{I}}(t)$, then the total number of paths of length 2 beginning at $v$ is at most $nd$, since $\deg(v)\leq n$. As there are $m$ vertices in $\mathsf{highdeg}(G)$, the number of paths in ${\mathcal{Q}}$ for this case is at most $mnd$. Summing both together leads to the claimed bound.
@@
($\operatorname{\mathbb{P}}\left(\{|{\mathcal{J}}_{2}|\geq 1\}\cap{\mathcal{A}}\cap{\mathcal{B}}_{t}|{\mathcal{F}}_{t}\right)$ $\leq\operatorname{\mathbb{P}}\left(\exists P\in{\mathcal{Q}}:\mathsf{weight}(P)\leq\delta|{\mathcal{F}}_{t}\right)\leq\sum_{P\in{\mathcal{Q}}}\operatorname{\mathbb{P}}(\mathsf{weight}(P)\leq\delta|{\mathcal{F}}_{t})$
$\stackrel{{\scriptstyle(a)}}{{\leq}}|{\mathcal{Q}}|\lambda^{2}\delta^{2}\stackrel{{\scriptstyle(b)}}{{\leq}}n^{1+o(1)}\delta^{2}=n^{-(2\beta-1)+o(1)}.$
@@
($\operatorname{\mathbb{P}}\left(\{|{\mathcal{J}}_{2}|\geq 1\}\cap{\mathcal{A}}|{\mathcal{F}}_{t}\right)$ ($\leq\operatorname{\mathbb{P}}\left(\{|{\mathcal{J}}_{2}|\geq 1\}\cap{\mathcal{A}}\cap{\mathcal{B}}_{t}|{\mathcal{F}}_{t}\right)+\operatorname{\mathbb{P}}({\mathcal{A}}\cap{\mathcal{B}}_{t}^{c}|{\mathcal{F}}_{t})$
@@
Lemma 5.1 now follows readily.
@@
By the definition of ${\mathcal{J}}_{1}$ and ${\mathcal{J}}_{2}$, we have that $I[t,t+n^{-\beta}]=|{\mathcal{J}}_{1}|$ if ${\mathcal{J}}_{2}=\emptyset$. It follows that
@@
($\operatorname{\mathbb{P}}\left(\left\{\left|I\left[t,t+\frac{1}{n^{\beta}}\right]-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}(t))\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right)\\ \leq\operatorname{\mathbb{P}}\left(\left\{\left||{\mathcal{J}}_{1}|-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}(t))\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right)+\operatorname{\mathbb{P}}(\{|{\mathcal{J}}_{2}|\geq 1\}\cap{\mathcal{A}})\leq n^{-(\beta-1/2)+o(1)}.$
5.2 Changes in the cut: Proof of Lemma 5.2
As a shorthand, denote $t_{1}:=t-n^{-\beta}$ and $t_{2}:=t-n^{-5}$. Consider the event
:=${\mathcal{B}}:=\{\forall v\in\mathsf{highdeg}(G),T(v)\notin(t_{1},t_{2}]\}.$
On this event, all vertices in ${\mathcal{I}}(t_{2})\setminus{\mathcal{I}}(t_{1})$ have degree at most $d$, hence
$|\operatorname{\mathsf{cut}}({\mathcal{I}}(t_{2}))-\operatorname{\mathsf{cut}}({\mathcal{I}}(t_{1}))|\leq\sum_{v\in{\mathcal{I}}(t_{2})\setminus{\mathcal{I}}(t_{1})}\deg(v)\leq d|{\mathcal{I}}(t_{2})\setminus{\mathcal{I}}(t_{1})|.$
Above, the first inequality follows since the size of the cut changes by at most $\deg(v)$ vertices when $v$ is added to the cascade.
We proceed by studying $|{\mathcal{I}}(t_{2})\setminus{\mathcal{I}}(t_{1})|$. We have that
@@
$\operatorname{\mathbb{E}}\left[|{\mathcal{I}}(t_{2})\setminus{\mathcal{I}}(t_{1})|\mathbf{1}({\mathcal{B}})|{\mathcal{F}}_{t_{1}}\right]$ =$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{v\in V\setminus{\mathcal{I}}(t_{1}):\deg(v)\leq d}\operatorname{\mathbb{P}}\left(\{T(v)\in(t_{1},t)\}\cap{\mathcal{B}}|{\mathcal{F}}_{t_{1}}\right)$
$\leq\sum_{v\in V\setminus{\mathcal{I}}(t_{1}):\deg(v)\leq d}\operatorname{\mathbb{P}}\left(T(v)\in(t_{1},t_{1}+n^{-\beta})|{\mathcal{F}}_{t_{1}}\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{v\in V\setminus{\mathcal{I}}(t_{1}):\deg(v)\leq d}\frac{d}{n^{\beta}}\stackrel{{\scriptstyle(c)}}{{=}}n^{1-\beta+o(1)}.$ (5.8)
Above, $(a)$ follows since all vertices in ${\mathcal{I}}(t_{2})\setminus{\mathcal{I}}(t_{1})$ have degree at most $d$ on ${\mathcal{B}}$, $(b)$ is due to Lemma 3.3 and $(c)$ follows since $d=n^{o(1)}$. Importantly, it follows immediately from (5.8) that the unconditional expectation can be bounded as $\operatorname{\mathbb{E}}[|{\mathcal{I}}(t_{2})\setminus{\mathcal{I}}(t_{1})|\mathbf{1}({\mathcal{B}})]\leq n^{1-\beta+o(1)}$. Putting everything together, we have that
@@
($\operatorname{\mathbb{P}}\left(\left\{|\operatorname{\mathsf{cut}}({\mathcal{I}}(t_{2}))-\operatorname{\mathsf{cut}}({\mathcal{I}}(t_{1}))|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\cap{\mathcal{B}}\right)$ ($\leq\operatorname{\mathbb{P}}\left(\left\{|{\mathcal{I}}(t_{2})\setminus{\mathcal{I}}(t_{1})|\geq\frac{n^{\gamma}}{d}\right\}\cap{\mathcal{B}}\right)$
@@
$\leq\frac{d}{n^{\gamma}}\operatorname{\mathbb{E}}\left[|{\mathcal{I}}(t_{2})\setminus{\mathcal{I}}(t_{1})|\mathbf{1}({\mathcal{B}})\right]$
$=n^{1-\beta-\gamma+o(1)}$ (5.9)
$\leq n^{-(2\beta-1)+o(1)},$ (5.10)
where the final inequality uses our assumption that $\gamma>\beta$. It remains to bound $\operatorname{\mathbb{P}}({\mathcal{A}}\cap{\mathcal{B}}^{c})$, which can be done using Lemma 3.8.
$\operatorname{\mathbb{P}}({\mathcal{A}}\cap{\mathcal{B}}^{c})\leq\sum_{v\in\mathsf{highdeg}(G)}\operatorname{\mathbb{P}}({\mathcal{A}}\cap\{T(v)\in(t_{1},t_{2})\})\leq mn^{-\left(\beta-\frac{1}{2}\right)+o(1)}=n^{-\left(\beta-\frac{1}{2}\right)+o(1)},$ (5.11)
where the final expression uses $m=n^{o(1)}$. The lemma now follows from (5.9) and (5.11). ∎
6 Lower bounds for $\alpha\in(0,1/2)$: Proof of Theorem 1.4